OpenStudy (anonymous):

For a cone, the formula r= the square root of 3V/3.14(Pi)H describes the relationship between the radius r of the base, volume V, and the height h. Find the volume if the radius is 4 inches and the cone is 10 inches high. (Round to the nearest tenth). Show work and how you got the answer please.

OpenStudy (valpey):

\[r = \sqrt{3V/(\pi*h)}\] \[r^2 = 3V/(\pi*h)\] \[1/3*(\pi*h)*r^2 = V\]

OpenStudy (anonymous):

r=sqrt(3V/piH) r^2=3V/piH after squaring both sides r^2(piH)=3V after multiplying both sides by piH r^2(piH)/3=V when r=4 and h=10 160pi/3=V

OpenStudy (valpey):

\[1/3*(\pi*10)*4^2 = 160\pi/3 \]

OpenStudy (anonymous):

53.3(pi)=V to the nearest tenth

OpenStudy (anonymous):

167.6

OpenStudy (anonymous):

Thankyou!

OpenStudy (anonymous):

welcome